OpenStudy (anonymous):

evaluate log2 16 the 2 is small...

OpenStudy (anonymous):

\[\log_{2} 16\]

OpenStudy (anonymous):

your job is to find \(y\) if \(2^y=16\)

OpenStudy (anonymous):

y=4?

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\] and so \[\log_2(16)=y\iff 2^y=16\] in other words your job is to find what number you would raise 2 to in order to get 16 yes it is 4 because \(2^4=16\)

OpenStudy (anonymous):

so \[\log_{5} 125\] is y=3...?

OpenStudy (anonymous):

you got it~

OpenStudy (anonymous):

so \(\log_5(125)=3\) logs are exponents, if you can switch back and forth between logarithmic and exponential forms quickly this will not be too hard

OpenStudy (anonymous):

gotcha!! now what about ln\[1/e^{5} \]

OpenStudy (anonymous):

ln in front of the problem